# Radial Basis Operator Networks

### v1

@@ 1.1 Background
The Laplace Neural Operator (LNO) [12] has recently become the new standard for operator networks in benchmarking performance. LNO was motivated as an extension to FNO in order to address primary limitations of the FNO, particularly its approximations for transient responses and non-periodic signals. To address this issue, LNO applies the pole-residue method to represent both the transient and steady-state response in the Laplace domain. This translates to better test performance on out-of-distribution (OOD) data in almost all contexts with few exceptions to include a reduced training cost and simpler network architecture. For this reason we have selected LNO as the primary comparison for our new operator network, including a problem scenario from [12] with small OOD error for the LNO predictions.
@@ 2 Methodology
The RBON is a numerical representation, $G_{\dagger}$, for an operator, $G:\mathcal{U}\to\mathcal{V}$, where $\mathcal{U}$ and $\mathcal{V}$ are infinite dimensional spaces, using radial basis functions. Following the work as shown in [13], we present, without proof, the universal approximation theorem for such a representation as well as extending the theorem to include NRBONs. The subsequent section details the precise implementation used for the experimental results.
@@ 2.1 Theoretical Foundation
Figure 1: The upper network is the branch net that receives the function input. The lower trunk net receives the sensor location input, and $\mathcal{L}$ denotes a linear transformation.
@@ 2.2 Practical Implementation
Initially, function inputs $u({\bm{x}})$ are sampled at $m$ point locations as defined in Section 2.1, while sensor location inputs ${\bm{y}}$ are processed directly. For both the trunk and branch networks, we select Gaussian functions for the RBF transformations on the input data. To determine the centers and spreads of the RBFs, we employ K-means clustering [14, 15] on the input data for each sub-network. The centers obtained from this clustering serve as the RBF centers, and the spread is calculated based on inter-cluster distances.
For weight parameter calculation, we compute the product of the RBF activations for the nodes in both the trunk and branch net. The weights, $\xi_{i}^{k}$, are then determined using the method of least squares. This is done for the branch output of every input $u^{m}_{j}$ for $j$ in the training index, iterated across the trunk output for each query location, ${\bm{y}}$. This process produces different values for the weight parameters for each ${\bm{y}}$ and the final weights are obtained by averaging across all query locations. A linear transform is applied to the final output, the weights of which are solved for directly.
@@
Concluding the description of the practical implementation, we note that solving for the network weights tends to yield more accurate predictions when using the Moore-Penrose inverse [16, 17] instead of a factorisation scheme. Using a least-mean-squares iterative approach also works, but results in weights that on average produce larger error in their predictions.
@@ 3.1.3 Euler-Bernoulli Beam Equation
Out ${\bf 8.6E{-3}(1.7E{-4})}$ ${\bf 2.3E{-2}(5.5E{-3})}$ $6.6\text{E}{-2}(7.0\text{E}{-3})$
@@ 3.1.4 Results
The results from these experiments can be seen in Table 1, which displays the average $L^{2}$ error for each function in the in-distribution and out-of-distribution test sets. The $L^{2}$ relative error is computed as follows
@@
where $v^{true}$ is the numerical representation of the ground truth at all query points in the space-time domain, and $v^{pred}$ is the network prediction for the same locations.
The margin of error (MOE) displayed in parentheses was calculated by multiplying the point estimate’s standard error with a critical value associated at the 95% confidence level. However, the strict interpretation of the confidence interval is lost due to the lack of independence between observations. Nevertheless, the MOE provides a notion of the variability present in the average error. Each of the different RBON types outperforms the LNO on the majority of the problems, with the NRBON outperforming LNO in every scenario for both ID and OOD data.
@@ 3.2.1 Results
The results in Table 2 demonstrate the ability of a RBON to accurately predict both local monthly average temperatures and global average temperatures. We trained both RBON and NRBON networks on available historical data, omitting the most recent two or five years from the training set to test the models’ forecasting accuracy.
Based solely on monthly $\text{CO}_{2}$ measurements and the month encoding for querying the output temperature, the RBON maintains an $L^{2}$ relative error of less than $10\%$, with NRBON performing similarly. Figure 3 displays a comparison between the trained networks’ global temperature predictions and actual global temperature readings. The left graph shows results when holding out the most recent two years, while the right graph illustrates the outcome when holding out the most recent five years of data. Figure 4 provides the same visual comparison for local temperatures recorded at the observatory that also publishes the atmospheric $\text{CO}_{2}$ measurements [22]. Note that temperature data for the local set was only available through 2018.
@@
Table 2: Average relative $L^{2}$ test error on temperature data with branch and trunk widths in braces.
Network 2/5 yr Global temp Local temp
RBON 2 0.02 $\{5,12\}$ 0.07 $\{10,9\}$
5 0.02 $\{5,12\}$ 0.07 $\{10,9\}$
NRBON 2 0.14 $\{27,5\}$ 0.13 $\{40,2\}$
5 0.15 $\{27,5\}$ 0.13 $\{40,2\}$
F-RBON 2 0.02 $\{2,9\}$ 0.04 $\{17,10\}$
5 0.01 $\{2,12\}$ 0.13 $\{4,2\}$
Figure 4: Two (left) and five (right) year local temperature predictions based on $\text{CO}_{2}$ input. Forecast values are in the shaded region.
Testing revealed that increasing the width of the branch and trunk networks enhances the model’s flexibility to match highly variable and erratic behavior. However, given highly oscillatory data, the plain RBON can occasionally produce peaks and valleys that deviate too far from the data range when increasing model width. In contrast, the NRBON can increase its network size without generating extreme peaks. Consequently, the smaller RBONs used yield a more stable regression appearance, while the larger NRBON networks produce outputs that attempt to capture more of the random extreme values. This results in a slightly higher error ($\leq 0.17$) for the NRBON, but a shape that more closely resembles the true graph. Figures 5 and 6 display the models’ output over a larger time interval, providing a clearer picture of this phenomenon upon close inspection. This difference is particularly evident in the right graph of Figure 6, where the NRBON attempts to better match the variability in the peaks, while the RBON produces a more steady-state prediction when trained on the smaller set of historical data.
@@
Figure 6: Two (left) and five (right) year average local temperature predictions based on $\text{CO}_{2}$ input.
@@ 4 Discussion and Conclusion
The RBON and its variants offer a simple yet powerful network architecture with prediction capabilities that yield errors smaller than the current leading operator network. The network’s compact size provides opportunities for enhanced interpretability and reduced computational load, allowing for exact solutions of network parameters. The majority of variation in different training cycles stems from the location and scale parameters of the RBFs, primarily due to K-means’ tendency to converge on local extrema. A quick solution to this issue is to run the K-means algorithm multiple times. Furthermore, the RBON serves as an excellent tool for scientific computing, where recent advancements have only begun to explore the potential of operator networks in various fields. Finally, the RBON’s ability to train on both real and complex-valued inputs, combined with its other strengths, makes it a promising candidate for applications in signal processing and computer vision tasks.

### v2

@@ 1.1 Background
The Laplace Neural Operator (LNO) [12] has recently become a benchmark standard for operator networks due to its improved handling of transient responses and non-periodic signals, limitations inherent in the Fourier Neural Operator (FNO). LNO achieves this by leveraging the pole-residue method to represent both transient and steady-state responses in the Laplace domain, leading to better test performance on out-of-distribution (OOD) data in most contexts. Additionally, LNO boasts a reduced training cost and a simpler network architecture. For these reasons, we have selected LNO as the primary comparison for our new operator network, alongside FNO and DeepONet. To thoroughly evaluate performance, we include a problem scenario from [12] that highlights LNO’s small OOD error in predictions.
@@ 2 Methodology
The RBON is a numerical representation, $G_{\dagger}$, for an operator, $G:\mathcal{U}\to\mathcal{V}$, where $\mathcal{U}$ and $\mathcal{V}$ are infinite dimensional spaces, using radial basis functions. Following the work as shown in [*]ChenRadial, we present, without proof, the universal approximation theorem for such a representation as well as extending the theorem to include NRBONs. The subsequent section details the precise implementation used for the experimental results.
@@ 2.1 Theoretical Foundation
Figure 1: The upper network is the branch net that receives the function input, and the lower trunk net receives the query location input. The product between the trunk and branch networks is represented by the Kronecker product, and $\mathcal{L}$ denotes a linear transformation.
@@ 2.2 Practical Implementation
From Theorem 2.1, recall $u^{m}\in\mathbb{R}^{m}$ represents the numerical approximation of the function $u$ sampled at $m$ locations, $G^{\dagger}$ is the network approximation of the operator, $G$, mapping $u^{m}$ to the function $v$ at the query location ${\bm{y}}\in R^{d}.$ Then, given input functions $u^{m}_{j}$ for $j\in\{1,\dots,J\}$, and query locations ${\bm{y}}_{\ell}$ for $\ell\in\{1,\dots,L\}$ where $J$ and $L$ denote the number of training input functions and query points, respectively, we outline the process for finding the network parameters.
RBF transformations. In both the trunk and branch networks we employ Gaussian functions for the RBF transformations, defined as
$\phi(\mathbf{x},\mathbf{c},\sigma)=\exp\left(-\frac{||\mathbf{x}-\mathbf{c}||^{2}}{2\sigma^{2}}\right)$
where $\mathbf{c}$ and $\sigma$ are the RBF centers and spreads. The RBF centers are determined using K-means clustering [19, 20] on the input data for each sub-network, with the spreads calculated based on inter-cluster distances. The branch and trunk network transformations on an input pair $\{u^{m}_{j},{\bm{y}}_{\ell}\}$, with $M$ and $N$ RBFs, are represented by the vectors
$\mathbf{b}(u^{m}_{j})=$ $[\phi(u^{m}_{j},\mathbf{c}_{1}^{b},\sigma_{1}^{b}),...,\phi(u^{m}_{j},\mathbf{c}_{M}^{b},\sigma_{M}^{b})]^{T}\,\,$
$\mathbf{t}({\bm{y}}_{\ell})=$ $[\phi({\bm{y}}_{\ell},\mathbf{c}_{1}^{t},\sigma_{1}^{t}),...,\phi({\bm{y}}_{\ell},\mathbf{c}_{N}^{t},\sigma_{N}^{t})]^{T}$
where $\mathbf{c}_{i}^{b},\mathbf{c}_{k}^{t}$ are the RBF centers and $\sigma_{i}^{b},\sigma_{k}^{t}$ are spreads for the associated branch and trunk networks.
Weight parameter calculation. For each query location ${\bm{y}}_{\ell}$, we first compute
$\bm{\Phi}_{\ell}=[\mathbf{b}(u^{m}_{1})\otimes\mathbf{t}({\bm{y}}_{\ell}),...,\mathbf{b}(u^{m}_{J})\otimes\mathbf{t}({\bm{y}}_{\ell})]$
where $\otimes$ denotes the Kronecker product, making $\bm{\Phi}_{\ell}$ of dimension $NM\times J$. The weights $\bm{\xi}_{\ell}$ of dimension $NM\times 1$, are then determined by solving
$\bm{\xi}_{\ell}^{T}\bm{\Phi}_{\ell}=[v_{1}({\bm{y}}_{\ell}),\cdots,v_{J}({\bm{y}}_{\ell})]$
using the Moore-Penrose inverse [22, 21]. This process yields $L$ weight vectors $\bm{\xi}_{\ell}$, for each query point. The final weight vector $\bm{\xi}$ is obtained by element-wise averaging across the $L$ vectors $\bm{\xi}_{\ell}$. Given the input $u^{m}$, the network approximation for the associated output function $v$ at query point ${\bm{y}}$ is then
$G^{\dagger}(u^{m})({\bm{y}})=\mathcal{L}(\bm{\xi}^{T}[\mathbf{b}(u^{m})\otimes\mathbf{t}({\bm{y}}_{\ell})])$
where $\mathcal{L}$ denotes a linear transformation applied to the final output whose parameters are solved for directly using the training data.
NRBON modification. The NRBON differs from RBON in normalizing the products of the branch and trunk outputs by dividing each element of the vector $[\mathbf{b}(u^{m})\otimes\mathbf{t}({\bm{y}}_{\ell})]$ by the vector’s sum. This normalization adjusts the computation of $\bm{\Phi}_{\ell}$ by its column totals.
@@
Concluding the description of the practical implementation, we note that the network weights can be solved for using an iterative approach such as least-mean-squares, but results in weights that on average produce larger error in their predictions.
@@ 3.1.3 Euler-Bernoulli Beam Equation
Out ${\bf 8.6E{-3}(1.7E{-4})}$ $2.3{\text{E}}{-2}(5.5{\text{E}}{-3})$ $6.6\text{E}{-2}(7.0\text{E}{-3})$
@@
FNO In $9.9{\text{E}}{-4}(2.3{\text{E}}{-5})$ $9.3{\text{E}}{-3}(1.2{\text{E}}{-3})$ $4.0{\text{E}}{-3}(6.1{\text{E}}{-3})$
Out $1.1{\text{E}}{-1}(1.4{\text{E}}{-3})$ ${\bf 1.7{\text{E}}{-2}(7.0{\text{E}}{-6})}$ $1.5{\text{E}}{-3}(2.2{\text{E}}{-4})$
[…]
Out $4.9{\text{E}}{-2}(3.4{\text{E}}{-5})$ $9.9{\text{E}}{-1}(2.0{\text{E}}{-6})$ $2.5{\text{E}}{-1}(1.4{\text{E}}{-2})$
@@ 3.1.4 Results
The results from these experiments can be seen in Table 1, which displays the average $L^{2}$ error for each function in the ID and OOD test sets. The $L^{2}$ relative error is computed as follows
@@
where $v^{\text{true}}$ represents the ground truth values at all query points in the space-time domain, and $v^{\text{pred}}$ represents the network’s predictions at the same locations.
The margin of error (MOE), shown in parentheses, was computed by multiplying the standard error of the point estimate by the critical value for a 95
Across the majority of problems, the RBON variants outperform the LNO, with the NRBON achieving consistently superior performance on both ID and OOD data. Overall, RBON variants collectively tend to outperform other operator networks, with one exception. Notably, operator networks generally exhibit smaller errors for the Beam equation due to their ability to accurately represent linear operators. Networks that leverage global representations—such as FNO, LNO, and F-RBON (which trains on data with global representations)—tend to generalize better, while other networks overfit ID data and have significantly worse performance on OOD data. This difference is especially noticeable with the OOD input data for the Wave problem due to its highly oscillatory behavior.
DeepONet initially suffered from overfitting, resulting in poor OOD performance, but early stopping significantly improved its OOD errors, albeit at the cost of slightly worse ID errors. However, this improvement came at the expense of efficiency: DeepONet required significantly larger sub-networks, with over 10,000 products between trunk and branch outputs, compared to fewer than 200 products in the RBONs.
@@ 3.2.1 Results
The results in Table 2 highlight the effectiveness of RBON in accurately predicting both local monthly average temperatures and global average temperatures. To evaluate the forecasting accuracy, we trained RBON and NRBON networks on historical temperature data, withholding the most recent two or five years from the training set for testing. In addition to these models, we compared their performance against LNO, DeepONet, FNO, and LSTM. This comprehensive evaluation demonstrates the robustness of RBON across diverse benchmarks, including traditional time-series approaches such as LSTM [16] and as well as other operator networks.
Based solely on monthly $\text{CO}_{2}$ measurements and the month encoding for querying the output temperature, the RBON maintains an $L^{2}$ relative error of less than $10\%$, with NRBON performing similarly. Figure 3 displays a comparison between the trained RBON networks’ global temperature predictions and actual global temperature readings. The left graph shows results when holding out the most recent two years, while the right graph illustrates the outcome when holding out the most recent five years of data. Interestingly, several networks—including RBON, F-RBON, DeepONet and LSTM—performed similarly on the smoother global temperature data. However, performance on the more variable local temperatures at the observatory publishing the atmospheric $\text{CO}_{2}$ measurements [26] provided a more clear distinction as RBON outperformed other networks, which struggled to capture the finer-scale variations in the data. Figure 4 provides the visual comparison for local temperatures versus the predictions from the RBON variants. Note that temperature data for the local set was only available through 2018.
@@
Table 2: Average relative $L^{2}$ test error on local and global temperature data as predicted based on atmospheric $\text{CO}_{2}$ input data.
RBON NRBON F-RBON LNO LSTM FNO DeepONet
Global temp: 2 yr 0.02 0.14 0.02 0.96 0.02 0.31 0.01
                  5 yr 0.02 0.15 0.01 0.97 0.02 0.44 0.01
Local temp: 2 yr 0.07 0.13 0.04 0.94 0.35 0.18 0.15
                 5 yr 0.07 0.13 0.13 0.95 0.51 0.22 0.14
Figure 4: Two (top) and five (bottom) year local temperature predictions based on $\text{CO}_{2}$ input. Forecast values are in the shaded region.
Testing revealed that increasing the width of the branch and trunk networks enhances the model’s flexibility to match highly variable and erratic behavior. However, given highly oscillatory data, the plain RBON can occasionally produce peaks and valleys that deviate too far from the data range when increasing model width. In contrast, the NRBON can increase its network size without generating extreme peaks. Consequently, the smaller RBONs used yield a more stable regression appearance, while the larger NRBON networks produce outputs that attempt to capture more of the random extreme values. This results in a slightly higher error ($\leq 0.17$) for the NRBON, but a shape that more closely resembles the true graph. Figures displaying the models’ output over a larger time interval, providing a clearer picture of this phenomenon upon close inspection, can be viewed in the Appendix. This difference is particularly evident in the right graph of Figure 6, where the NRBON attempts to better match the variability in the peaks, while the RBON produces a more steady-state prediction when trained on the smaller set of historical data.
@@ 4 Discussion and Conclusion
The RBON and its variants offer a simple yet powerful network architecture with prediction capabilities that yield errors smaller than the current leading operator network. The network’s compact size provides opportunities for enhanced interpretability and reduced computational load, allowing for exact solutions of network parameters. Most variation across training cycles arises from the location and scale parameters of the RBFs, largely due to K-means’ tendency to converge on local extrema. This variability can lead to errors differing by several orders of magnitude between runs of the K-means algorithm. A practical solution is to run K-means multiple times and select the configuration that minimizes the overall within-cluster distances. Furthermore, the RBON serves as an excellent tool for scientific computing, where recent advancements have only begun to explore the potential of operator networks in various fields. Finally, the RBON’s ability to train on both real and complex-valued inputs, combined with its other strengths, makes it a promising candidate for applications in signal processing and computer vision tasks.
@@ A.1 Temperature model visuals
A.1.1 RBON and NRBON visuals
This section provides visuals for the temperature data and the NRBON and RBON predictions based on input $\text{CO}_{2}$ data on a larger time axis to provide a more complete view of the overarching trends. Specifically, Figures 5 and 6 provide the global and local temperature predictions over a much larger time period than the graphs present in Section 3.2.1.
Figure 5: Two (top) and five (bottom) year average global temperature predictions based on $\text{CO}_{2}$ input.
Figure 6: Two (top) and five (bottom) year average local temperature predictions based on $\text{CO}_{2}$ input.
A.1.2 F-RBON visuals